OpenStudy (anonymous):

If x^2-12x+c is a perfect square trinomial, what is the value of c?

OpenStudy (anonymous):

(b/2)^2

OpenStudy (anonymous):

Please explain:)

OpenStudy (ash2326):

x^2-2*6x+c=0 standard perfect square trinomial is the form of a^2-2a*b+b^2=0 here a=1 -2ab=-12 a=1 so b=6 c=b^2=36

OpenStudy (anonymous):

Thankyou!:)